# Simplified Mamba with Disentangled Dependency Encoding for Long-Term Time Series Forecasting

### v2

@@ 4.1 Why Mamba? A Deep Dive into Its Suitability for LTSF
Assumption 2. Patching enhances the semantic dependency of a sequence (Nie et al. (2022)).
@@
The experimental results in Table 2 show a significant improvement in the Transformer after patching compared to the linear model after patching, demonstrating its stronger semantic learning capability. The Linear model, due to its limited expressive power, struggles to effectively capture semantic dependency, further validating the rationale behind our experimental approach. In addition, compared to the Linear model, Mamba with the patching strategy also exhibits notable improvements and achieves the best performance, which verifies its strength in capturing semantic dependency.
In summary, our evaluation results above indicate that Mamba is the only model capable of simultaneously capturing order and semantic dependencies, making it particularly well-suited for LTSF, where both order and semantic dependencies are critical.
Table 2: Results of the linear model, Mamba, and Transformer w/ or w/o patching on ETTm1 dataset.
@@
4.2 The Pitfalls of Nonlinearity: Overfitting in Deep LTSF Models
@@
The success of Linear models has prompted researchers to reconsider the utility of more complex architectures, such as Transformers (Zeng et al. (2023)). Initially designed for NLP tasks, these models emphasize semantic dependency over order dependency and incorporate multiple nonlinear activation functions. Consequently, we hypothesize that these nonlinear activation functions may increase the risk of overfitting in time series data, which possess lower information density compared to natural language.
@@
To test this hypothesis, we carry out ablation studies using nonlinear activation functions across multiple backbone models: MLP, Transformer, and Mamba, along with their respective variants that incorporate a patching strategy. As shown in Table 3, including nonlinear activation functions negatively impacts the model’s performance. Interestingly, removing these functions results in notable performance improvements. Figure 1 further illustrates this trend, showing that MLP, Mamba, and Transformer exhibit varying levels of overfitting, with the nonlinear versions overfitting more than their linear counterparts. This observation implies that while nonlinear activation functions enhance a model’s capacity for semantic dependency learning, they may simultaneously impair its ability to generalize from temporal patterns in time series data.
@@
Table 3: Ablation study on nonlinear activation function on ETTm1 dataset, we report the average performance. ’Original’ means vanilla model, ’-n’ means removing the nonlinear activation function and ’Patch+’ means using patching.
  Models MLP Mamba Patch+Mamba Transformer Patch+Transformer
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
Original 0.415 0.421 0.674 0.598 0.402 0.412 0.815 0.682 0.414 0.422
Original-n 0.406 0.411 0.635 0.585 0.399 0.414 0.653 0.600 0.406 0.417
Improvement 2.17% 2.37% 5.79% 2.17% 0.75% -0.48% 19.88% 12.02% 1.93% 1.18%
@@
Our analysis further reveals that the adverse effects of nonlinear activation functions are most pronounced in Transformer architectures, with Mamba and MLP following. Notably, this order is inversely related to the model’s ability to capture order dependency. This suggests that a model’s ability to capture order dependency can partially mitigate the overfitting introduced by nonlinear activation functions, thereby enhancing its capacity to generalize temporal patterns.
Furthermore, after applying the patching strategy, we observe that removing nonlinear activation functions results in fewer performance variability in Mamba and Transformer models. The patching strategy increases the semantic density of the time series data, allowing nonlinear models to better utilize their capacity without overfitting to sparse information. However, despite the ability of patching to mitigate overfitting to some extent, removing nonlinear activation functions remains overall beneficial, leading to more stable and generalizable models.
Figure 1: Training curves of three models and their variants with nonlinear activation removed.
5 SAMBA
As aforementioned, adapting Mamba to LTSF presents two non-trivial challenges: over-reliance on semantic dependency and the difficulty in leveraging cross-variate dependency without compromising the individual characteristics of each variate. To address these challenges, we analyze the training dynamic of Mamba in Section 4.2 and identify an overfitting issue caused by nonlinear activation function. Inspired by this, we proposed SAMBA, illustrated in Figure 2, with a disentangled dependency encoding strategy to avoid mutual interference between time and variate dimensions.
Figure 2: The overall framework of SAMBA.
5.1 Simplifying Mamba for LTSF
@@
Although attention mechanism excel in learning semantic dependency, its inherent permutation invariance constrains its ability to accurately capture order dependency, despite the use of opsitional encoding (Zeng et al. (2023)). Compared to Transformer, Mamba’s strength in LTSF lies in its use of selective SSM that is similar to self-attention but processes time series recursively. This approach enables Mamba to capture order dependency and semantic dependency simultaneously, as demonstrated in Section 4.1. Formally, selective SSM can be described as:
@@
Based on discrete SSM that has been theoretically proven to capture long-range order dependency (Gu et al. (2021a)), Mamba introduces a selection mechanism that makes $\mathbf{B}$, $\mathbf{C}$ and $\bm{\Delta}$ data dependent, allowing it to capture semantic dependency as well:
@@
This branch is designed to efficiently capture both order and semantic dependencies. Based on our analysis in Section 4.2, we propose a SAMBA block that removes the nonlinear activation function between the Conv1D and SSM layers to mitigate the overfitting issue. Our rationale is that 1D convolution, similar to a linear layer, is sufficient to learn a powerful time series representation.
@@
Notice that we retain the nonlinear activation within the gating mechanism (right in Figure 3) to ensure the stability and robustness of the learning process ( Gu & Dao (2023); Chung et al. (2014)).
Figure 3: The framework of a SAMBA block.
5.2 Disentangled Dependency Encoding
Previous CD approaches encode cross-time dependency (order and semantic dependencies) and cross-variate dependency sequentially. However, this can introduce less relevant cross-time dependency from other variates, which negatively impact the informativeness of the embedding for the target variate. To address this, we propose a theoretically sound disentangled encoding strategy that processes cross-time and cross-variate dependencies in parallel. By aggregating information from both dimensions and mapping them into a unified space, our strategy ensures that these dependencies are fully leveraged while minimizing the negative interference between them.
@@ 5.2 Disentangled Dependency Encoding
Cross-time Dependency Modeling. To accurately capture cross-time dependency, we independently model each variate. First, we introduce the learnable additive position encoding $\mathbf{W}_{\mathrm{pos}}\in\mathbb{R}^{J\times D}$ to enhance the order information of the patches, modifying the embedding as $\mathbf{E}_{time}^{i}\leftarrow\mathbf{E}_{time}^{i}+\mathbf{W}_{\mathrm{pos}}$, where $\mathbf{E}_{time}^{i}\in\mathbb{R}^{J\times D}$. Subsequently, we process each channel independently through a single SAMBA block, equipped with residual connections and Layer Normalization (LN) (Ba et al. (2016)): $\mathbf{E}_{time}^{i}=\text{LN}(\text{{\sc SAMBA}}(\mathbf{E}_{time}^{i})+\mathbf{E}_{time}^{i})$. This process effectively extracts the order dependency and semantic dependency of each channel, producing a patch representation $\mathbf{E}_{time}\in\mathbb{R}^{N\times J\times D}$ enriched with temporal information.
Cross-variate Dependency Modeling. Existing methods that introduce cross-variate dependency often conflate temporal and variate information, reducing the distinguishability of each channel (Liu et al. (2023)). This can negatively impact model performance. To overcome this limitation, we design an adjustable bidirectional SAMBA module to capture complex cross-variate dependency in a disentangled way without blurring the distinction between the time and variate dimensions. The input sequence is $\mathbf{E}_{var}^{i}\in\mathbb{R}^{N\times D}$. Each input sequence through the module will generate a forward result $\mathbf{E}_{fo}^{i}$ and a backward result $\mathbf{E}_{ba}^{i}$, maintaining the disorder of time series in variate dimensions. Thereafter, we employ a flexible aggregation operation to integrate the forward and backward information from the variate dimensions of time series sequences and use layer normalization to increase the convergence: $\mathbf{E}_{var}^{i}=\text{LN}(\alpha\mathbf{E}_{fo}^{i}+\beta\mathbf{E}_{ba}^{i}+\mathbf{E}_{var}^{i})$ where $\alpha$ and $\beta$ are hyperparameters. This strategy produces patch representations $\mathbf{E}_{var}\in\mathbb{R}^{J\times N\times D}$ that encompass complex cross-variate dependency.
@@
Theoretical Discussions. Finally, we theoretically analyze the sufficient expressiveness of our disentanglement strategy compared to LSTF models that alternatively encode cross-time and cross-variate dependencies. To begin with, let $\Phi:\mathbb{R}^{J\times D}\rightarrow\mathbb{R}^{J\times D}$ and $\Psi:\mathbb{R}^{N\times D}\rightarrow\mathbb{R}^{N\times D}$ denote a cross-time encoder and a cross-variate encoder, respectively. With a little abuse of notation, a broadcasting mechanism will be applied when encoding 3D tensor inputs. Then our disentangled model can be denoted as $\mathcal{F}_{d}(\cdot)=\text{FFN}(\Phi(\cdot)||\Psi(\cdot)),$ which encodes the two types of dependencies in parallel. Based on Theorem 3.5 in Gao & Ribeiro (2022) and the evidence from Liu et al. (2023), we focus on comparing our model with time-then-variate LSTF models, which can be defined as $\mathcal{F}_{ttv}(\cdot)=\text{FFN}(\Psi(\Phi(\cdot))).$ Without loss of generality, we only consider the single-layer setting. We further assume that the cross-time encoder $\Phi(\cdot)$ can be decomposed into $\Phi(\cdot)=\phi(\mathcal{O}(\cdot),\mathcal{S}(\cdot))+\mathcal{Z}(\cdot),$ where $\mathcal{O}(\cdot)$ and $\mathcal{S}(\cdot)$ extract order and semantic dependency respectively, $\mathcal{Z}(\cdot)$ is the noisy component, and $\phi(\cdot)$ is a mapping. Let $\mathbf{c}^{i}:=\mathbf{c}^{i}_{T+1:T+S}$ for simplicity. Our main theorem is as follows.
@@
For any variate $i$, $1\leq i\leq N,$ if (1) $I(\mathbf{c}^{i};\mathcal{Z}(\mathbf{E})^{j}|\mathcal{O}(\mathbf{E})^{j},\mathcal{S}(\mathbf{E})^{j})=0,\forall j,1\leq j\leq N$ and (2) the dependencies of different variates satisfy $I(\mathbf{c}^{i};\mathcal{S}(\mathbf{E})^{-i}|\Phi(\mathbf{E})^{i},\mathcal{O}(\mathbf{E})^{-i})\leq\epsilon$ for some $\epsilon\geq 0,$ where $\mathcal{O}(\mathbf{E})^{-i}$($\mathcal{S}(\mathbf{E})^{-i}$) denotes the order(semantic) dependencies of variates except from the $i$-th one, then the informativeness of the representations output by the disentangled model and the time-then-variate model satisfy $H(\mathbf{c}^{i}|\mathcal{F}_{d}(\mathbf{E})^{i})\leq H(\mathbf{c}^{i}|\mathcal{F}_{ttv}(\mathbf{E})^{i})+\epsilon.$

### v3

@@ 4.1. Why State Space model? A Deep Dive into Its Suitability for LTSF
Assumption 2. Patching enhances the semantic dependency of a sequence (Nie et al., 2022).
@@
The experimental results presented in Table 1 underscore a substantial performance gain in the Transformer after applying the patching strategy, significantly outperforming the Linear model under the same conditions. This highlights the Transformer’s superior capability in semantic learning. The Linear model, constrained by its limited expressive power, fails to effectively capture complex semantic dependencies, further justifying the rationale behind our experimental design. Moreover, Mamba, when integrated with the patching strategy, demonstrates even greater improvements over the Linear model, achieving the best overall performance. This result strongly validates Mamba’s effectiveness in modeling semantic dependencies. Notably, as SegRNN has already established the efficacy of the patching strategy, we focus solely on state space models with Mamba in this analysis.
To summarize, the evaluation results presented above demonstrate that SSMs, including SegRNN and Mamba, uniquely possess the ability to capture both order and semantic dependencies. This capability makes them particularly suitable for LTSF, where maintaining both order and semantic relationships is crucial.
Table 1. Results of the linear model, Mamba, and Transformer w/ or w/o patching on ETTm1 dataset.
@@
4.2. The Pitfalls of Nonlinearity: Overfitting in Deep LTSF Models
Figure 1. Sharpness at the end of the training. Transformer-n refers to removing the activation functions in Transformer, and Mamba-n indicates that the Mamba removes activation functions.
@@ 4.2. The Pitfalls of Nonlinearity: Overfitting in Deep LTSF Models
The success of Linear models has prompted researchers to reconsider the utility of more complex architectures, such as Transformers (Zeng et al., 2023). Initially designed for NLP tasks, these models emphasize semantic dependency over order dependency and incorporate multiple nonlinear activation functions. Consequently, we hypothesize that these nonlinear activation functions may increase the risk of overfitting in time series data, which possess lower information density compared to natural language.
@@
To test this hypothesis, we carry out ablation studies using nonlinear activation functions across multiple backbone models: Transformer, and Mamba, along with their respective variants that incorporate a patching strategy. As shown in Table 2, including nonlinear activation functions negatively impacts the model’s performance. Interestingly, removing these functions results in notable performance improvements. Figure 1 further illustrates this trend by showing that sharpness—reflecting the curvature of the loss surface—is typically associated with increased overfitting. Notably, removing the nonlinear activation function leads to a reduction in sharpness, thereby indicating a mitigation of the overfitting problem. This observation implies that while nonlinear activation functions enhance a model’s capacity for semantic dependency learning, they may simultaneously impair its ability to generalize from temporal patterns in time series data.
@@
Table 2. Ablation study on nonlinear activation function on ETTm1 dataset. ’Original’ means vanilla model, ’-n’ means removing the nonlinear activation function and ’Patch+’ means using patching.
  Models Mamba Patch+Mamba Transformer Patch+Transformer
Metric MSE MAE MSE MAE MSE MAE MSE MAE
Original 0.674 0.598 0.395 0.419 0.815 0.682 0.414 0.422
Original-n 0.635 0.585 0.391 0.403 0.653 0.600 0.406 0.417
Improvement 5.79% 2.17% 1.01% 1.41% 19.88% 12.02% 1.93% 1.18%
Table 3. Order dependency analysis. Based on Assumption 1, we compare each model’s performance variation before and after randomly shuffling the temporal order of testing data points. O.MSE and O.MAE are evaluated in the original test set. S.MSE and S.MAE are evaluated in the shuffling test set.Avg.Drop means the average performance degradation.
Models
Linear Model
Mamba
SegRNN
Transformer
iTransformer
Metric
O.MSE
S.MSE
O.MAE
S.MAE
O.MSE
S.MSE
O.MAE
S.MAE
O.MSE
S.MSE
O.MAE
S.MAE
O.MSE
S.MSE
O.MAE
S.MAE
O.MSE
S.MSE
O.MAE
S.MAE
ETTm1
96
0.383
0.988
0.400
0.697
0.517
0.922
0.508
0.688
0.336
0.982
0.375
0.659
0.643
0.884
0.575
0.643
0.345
0.892
0.378
0.610
192
0.413
0.986
0.415
0.697
0.575
0.931
0.546
0.699
0.376
0.979
0.396
0.662
0.805
1.01
0.664
0.730
0.383
0.903
0.395
0.617
336
0.441
0.987
0.435
0.698
0.730
0.957
0.634
0.703
0.402
1.070
0.415
0.699
0.882
1.12
0.737
0.817
0.423
0.923
0.420
0.630
720
0.497
0.992
0.469
0.704
0.873
0.973
0.704
0.723
0.458
1.169
0.445
0.739
0.928
1.12
0.752
0.800
0.489
0.932
0.456
0.641
Avg. Drop
127.97%
62.55%
40.37%
17.60%
167.18%
69.16%
22.40%
6.55%
122.56%
51.5%
Exchange
96
0.0832
0.210
0.201
0.332
1.260
1.401
0.915
0.943
0.083
0.252
0.200
0.356
0.730
0.738
0.782
0.722
0.0869
0.242
0.207
0.358
192
0.179
0.325
0.299
0.414
1.398
1.626
1.040
1.060
0.176
0.374
0.296
0.436
1.304
1.284
0.913
0.949
0.179
0.374
0.301
0.450
336
0.338
0.521
0.418
0.534
1.835
1.921
1.111
1.141
0.332
0.531
0.416
0.537
1.860
1.862
1.090
1.085
0.331
0.555
0.417
0.557
720
0.903
1.167
0.714
0.822
3.940
4.023
1.687
1.697
0.935
1.190
0.736
0.836
3.860
3.865
1.684
1.685
0.856
1.202
0.698
0.841
Avg. Drop
47.89%
28.80%
6.38%
1.85%
53.80%
31.37%
-0.06%
-0.63%
63.33%
35.89%
Traffic
96
0.656
1.679
0.403
0.882
0.634
1.859
0.363
0.967
0.651
1.634
0.327
0.850
0.788
1.846
0.467
0.918
0.396
1.945
0.271
0.976
192
0.609
1.665
0.382
0.876
0.637
2.387
0.375
1.120
0.657
1.633
0.332
0.851
0.850
1.715
0.483
0.878
0.416
1.751
0.279
0.916
336
0.615
1.671
0.385
0.880
0.674
2.428
0.386
1.127
0.678
1.636
0.341
0.853
0.836
1.676
0.480
0.861
0.430
1.801
0.287
0.938
720
0.656
1.697
0.405
0.882
0.732
1.751
0.414
0.916
0.676
1.779
0.327
0.900
0.859
1.746
0.490
0.880
0.559
1.978
0.375
0.971
Avg. Drop
164.67%
123.49%
214.72%
168.53%
149.33%
160.29%
109.51%
84.22%
315.05%
213.61%
@@
Our analysis further reveals that the adverse effects of nonlinear activation functions are most pronounced in Transformer-based architectures, with Mamba trailing. Interestingly, this order is inversely correlated with the model’s ability to capture order dependency. This implies that a model’s capability to grasp order dependency can partially mitigate the overfitting caused by nonlinear activation functions, thereby enhancing its potential to generalize temporal patterns.
5. SDE: Simplified and Disentangled Dependency Encoding
As aforementioned, adapting SSMs to LTSF presents two non-trivial challenges. First, existing SSMs tend to overly rely on semantic dependencies, which may obscure the intrinsic temporal dynamics. Second, effectively capturing cross-variate dependencies without compromising the distinct characteristics of individual variates. To address these challenges, we propose SDE (Simplified and Disentangled Dependency Encoding), which consists of two core components, each designed to mitigate one of these issues.
The first component, Simplification Mechanism, is motivated by our analysis of Mamba’s training dynamics in Section 4.2, where we identify an overfitting issue stemming from excessive nonlinear activation. Inspired by this observation, we introduce a structured simplification strategy that removes unnecessary nonlinear operations traditionally used for learning complex representations.
The second component is a theoretically grounded disentangled dependency encoding strategy, which effectively disentangles temporal (order and semantic dependencies) and cross-variate dependencies. Unlike conventional approaches that encode these dependencies sequentially—often leading to entangled representations that dilute meaningful relationships—our framework explicitly models them in parallel. This disentanglement ensures that cross-variate interactions are captured effectively while preserving the distinct temporal dynamics of each variate.
To facilitate a clear and structured introduction to our approach, we take Mamba as a representative example to illustrate the details and procedural steps of the SDE framework. The resulting architecture, SDE-Mamba, is illustrated in Figure 3.
Figure 2. The overall framework of SDE-Mamba.
Figure 3. A S-Mamba block.
5.1. Simplification Mechanism for Mamba
@@
Although attention mechanism excel in learning semantic dependency, its inherent permutation invariance constrains its ability to accurately capture order dependency, despite the use of psitional encoding (Zeng et al., 2023). Compared to Transformer, Mamba’s strength in LTSF lies in its use of selective SSM that is similar to self-attention but processes time series recursively. This approach enables Mamba to capture order dependency and semantic dependency simultaneously, as demonstrated in Section 4.1. Formally, selective SSM can be described as:
@@
Based on discrete SSM that has been theoretically proven to capture long-range order dependency (Gu et al., 2021), Mamba introduces a selection mechanism that makes $\mathbf{B}$, $\mathbf{C}$ and $\bm{\Delta}$ data dependent, allowing it to capture semantic dependency as well:
@@
This branch is designed to efficiently capture both order and semantic dependencies. Building on our analysis in Section 4.2 and the proposed Simplification Mechanism, we eliminate the nonlinear activation function between the Conv1D and SSM layers in Mamba, leading to the development of S-Mamba. This design choice effectively mitigates the overfitting issue, as 1D convolution, similar to a linear layer, is sufficiently expressive to capture complex time series representations without introducing unnecessary nonlinearity.
@@
Notice that we retain the nonlinear activation within the gating mechanism (right in Figure 3) to ensure the stability and robustness of the learning process (Gu and Dao, 2023; Chung et al., 2014).
5.2. Disentangled Dependency Encoding
Previous CD approaches encode cross-time dependency (order and semantic dependencies) and cross-variate dependency sequentially. However, this paradigm can introduce less relevant cross-time dependency from other variates, negatively impacting the informativeness of the embedding for the target variate. To address this, we propose a theoretically sound disentangled encoding strategy that processes cross-time and cross-variate dependencies in parallel. By aggregating information from both dimensions and mapping them into a unified space, our strategy ensures that these dependencies are fully leveraged while minimizing the negative interference between them.
@@ 5.2. Disentangled Dependency Encoding
Cross-time Dependency Modeling. To accurately capture cross-time dependency, we independently model each variate. First, we introduce the learnable additive position encoding $\mathbf{W}_{\mathrm{pos}}\in\mathbb{R}^{J\times D}$ to enhance the order information of the patches, modifying the embedding as $\mathbf{E}_{\text{time}}^{i}\leftarrow\mathbf{E}_{\text{time}}^{i}+\mathbf{W}_{\mathrm{pos}}$, where $\mathbf{E}_{\text{time}}^{i}\in\mathbb{R}^{J\times D}$. Subsequently, we process each channel independently through a single $\mathcal{F}$ block, equipped with residual connections and Layer Normalization (LN) (Ba et al., 2016), formulated as $\mathbf{E}_{\text{time}}^{i}=\text{LN}(\mathcal{F}(\mathbf{E}_{\text{time}}^{i})+\mathbf{E}_{\text{time}}^{i})$. Here, $\mathcal{F}$ represents a generic sequence encoder, which can be instantiated as a Transformer, Mamba, or RNN, depending on the specific modeling needs. In our present implementation, we utilize S-Mamba as the encoder. This process effectively extracts both order and semantic dependencies for each variate, producing a patch representation $\mathbf{E}_{\text{time}}\in\mathbb{R}^{N\times J\times D}$ enriched with temporal information.
Cross-variate Dependency Modeling. Existing approaches to modeling cross-variate dependency often entangle temporal and variate information, thereby blurring the distinction between individual channels (Liu et al., 2023). This entanglement can degrade model performance by introducing spurious correlations and reducing the informativeness of learned representations. To overcome this limitation, we propose a unified cross-variate dependency encoder within our disentangled dependency encoding strategy, designed to disentangle these dependencies and preserve the structural integrity of each variate. In particular, this encoder is a modular component that can be instantiated with various architectures, such as Mamba, Transformers, or RNNs, allowing for adaptability to different modeling paradigms.
In our current instantiation, we employ S-Mamba as the cross-variate dependency encoder. SSMs, including Mamba and RNNs, inherently process inputs sequentially, which is not naturally suited for capturing unordered cross-variate dependency. To address this, we adapt S-Mamba to operate on an unordered set of variates by simply applying it to a random permutation of the variate order.
Formally, given an input sequence $\mathbf{E}_{\text{var}}^{i}\in\mathbb{R}^{N\times D}$, we apply S-Mamba over a randomly permuted variate order, generating the transformed representation $\mathbf{E}_{\text{fo}}^{i}$. To effectively integrate cross-variate dependency while preserving the distinct characteristics of each variate, we independently process patches corresponding to the same time step along the variate dimension, formulated as $\mathbf{E}_{\text{var}}^{i}=\text{LN}(\mathbf{E}_{\text{fo}}^{i}+\mathbf{E}_{\text{var}}^{i})$. This encoding strategy produces patch representations $\mathbf{E}_{\text{var}}\in\mathbb{R}^{J\times N\times D}$, effectively capturing complex cross-variate dependencies while maintaining a disentangled and structured representation.
@@
Theoretical Discussions. Finally, we theoretically analyze the sufficient expressiveness of our disentanglement strategy compared to LSTF models that alternatively encode cross-time and cross-variate dependencies. To begin with, let $\Phi:\mathbb{R}^{J\times D}\rightarrow\mathbb{R}^{J\times D}$ and $\Psi:\mathbb{R}^{N\times D}\rightarrow\mathbb{R}^{N\times D}$ denote a cross-time encoder and a cross-variate encoder, respectively. With a little abuse of notation, a broadcasting mechanism will be applied when encoding 3D tensor inputs. Then our disentangled model can be denoted as $\mathcal{F}_{d}(\cdot)=\text{FFN}(\Phi(\cdot)||\Psi(\cdot)),$ which encodes the two types of dependencies in parallel. Based on Theorem 3.5 in (Gao and Ribeiro, 2022) and the evidence from (Liu et al., 2023), we focus on comparing our model with time-then-variate LSTF models, which can be defined as $\mathcal{F}_{ttv}(\cdot)=\text{FFN}(\Psi(\Phi(\cdot))).$ Without loss of generality, we only consider the single-layer setting. We further assume that the cross-time encoder $\Phi(\cdot)$ can be decomposed into $\Phi(\cdot)=\phi(\mathcal{O}(\cdot),\mathcal{S}(\cdot))+\mathcal{Z}(\cdot),$ where $\mathcal{O}(\cdot)$ and $\mathcal{S}(\cdot)$ extract order and semantic dependency respectively, $\mathcal{Z}(\cdot)$ is the noisy component, and $\phi(\cdot)$ is a mapping. Let $\mathbf{c}^{i}:=\mathbf{c}^{i}_{T+1:T+S}$ for simplicity. Our main theorem is as follows.
@@
For any variate $i$, $1\leq i\leq N,$ if (1) $I(\mathbf{c}^{i};\mathcal{Z}(\mathbf{E})^{j}|$ $\mathcal{O}(\mathbf{E})^{j},\mathcal{S}(\mathbf{E})^{j})=0,\forall 1\leq j\leq N$ and (2) the dependencies of different variates satisfy $I(\mathbf{c}^{i};\mathcal{S}(\mathbf{E})^{-i}|\Phi(\mathbf{E})^{i},\mathcal{O}(\mathbf{E})^{-i})\leq\epsilon$ for some $\epsilon\geq 0,$ where $\mathcal{O}(\mathbf{E})^{-i}$($\mathcal{S}(\mathbf{E})^{-i}$) denotes the order(semantic) dependencies of variates except from the $i$-th one, then the informativeness of the representations output by the disentangled model and the time-then-variate model satisfy $H(\mathbf{c}^{i}|\mathcal{F}_{d}(\mathbf{E})^{i})\leq H(\mathbf{c}^{i}|\mathcal{F}_{ttv}(\mathbf{E})^{i})+\epsilon.$